OpenStudy (anonymous):

FAN AND MEDAL Ben has a 40% chance of being selected as the president of the school sports club and a 70% chance of being elected editor of the school magazine. What is the probability that Ben will be selected as the president of the sports club and will also be elected editor of the school magazine? 0.3% 1.1% 14% 28%

OpenStudy (mathmate):

When the success or failure of an event A does not affect the outcome of another even B, then A and B are independent. The probability of both events happening, P(A∩B)=P(A)*P(B). If the selection of the president and the editor is independent, then the probability of getting selected for both posts is the product of the probabilities of being selected for either post.

OpenStudy (anonymous):

ohh.....

OpenStudy (anonymous):

@starlord6200 do u know?

OpenStudy (anonymous):

@jordanjamesbay @tylermcmullen23

OpenStudy (anonymous):

what do u think

OpenStudy (anonymous):

idk? i don't it is A and B but idk

OpenStudy (anonymous):

hmm idk eather sorry

OpenStudy (anonymous):

its okay

OpenStudy (anonymous):

sorry i didn't mean to close it!

OpenStudy (mathmate):

@kiamousekia If you had time to read both paragraphs that I posted earlier, you should have a good chance of getting the answer by yourself.

OpenStudy (anonymous):

fine

OpenStudy (anonymous):

so did u figure it out yet

OpenStudy (anonymous):

no i got it wrong but it is fine

OpenStudy (mathmate):

... and you do not worry about why you got it wrong? :(

OpenStudy (anonymous):

i said C and it was wrong

OpenStudy (mathmate):

Was C a guess? "If the selection of the president and the editor is independent, then the probability of getting selected for both posts is the product of the probabilities of being selected for either post." Note: product = result of a multiplication, example: product of 2 and 3 = 2*3=6 P(president)=0.4 P(editor)=0.7 P(president and editor)=0.4*0.7=0.28

OpenStudy (anonymous):

yeah 14%!